OpenStudy (anonymous):

Tell whether the lines for the pair of equations are parallel, perpandicular, or neither. Y=-2/3x+1 2x-3y=-3

OpenStudy (anonymous):

-3y = -2x-3 y = 2/3x+1 y= -2/3x+1 y = mx+b m is the slope so parallel when the two equations have the same slope perpendicular when the slope of the equations are reciprocal with one another and have different sign

OpenStudy (anonymous):

So perpendicular? :) @helpme1.2

OpenStudy (anonymous):

example of perpendicular slope; if the first slope was 5/6 its perpendicular would be -6/5

OpenStudy (anonymous):

what do you notice about the 2 slopes for your problem

OpenStudy (anonymous):

they are the same! so they are parallel!? @helpme1.2

OpenStudy (anonymous):

one of them is negative and the one one is positive -2/3 not equal to 2/3

OpenStudy (anonymous):

so they are not the same meaning they are perpendicular? @helpme1.2

OpenStudy (anonymous):

-2/3 perpendicular would be 3/2 you forgot the other option Neither